Bibliography

Bruning-Guillemin1

[1] J. Brüning and V.W. Guillemin (Editors), Fourier integral operators, Springer-Verlag, Berlin,

Heidelberg, New York, Tokyo, 1994.

Brylinski-Getzler1

[2] J.-L. Brylinski and E. Getzler, The homology of algebras of pseudo-differential symbols and

the noncommutative residue, K-theory 1 (1987), 385–403.

MR86h:46002

[3] F. G. Friedlander, Introduction to the theory of distributions, Cambridge University Press,

Cambridge, 1982.

Friedlander2

[4] F.G. Friedlander, Introduction to the theory of distributions, Cambridge University Press,

Cambridge, London, New York, Melbourne, ?

Guillemin2

[5] V.W. Guillemin, A new proof of Weyl’s formula on the asymptotic distribution of eigenvalues,

Adv. Math. 55 (1985), 131–160.

Guillemin3

[6] , Residue traces for certain algebras of Fourier integral operators, J. Funct. Anal. 115

(1993), 391–417.

Hormander1

[7] L. Hörmander, Fourier integral operators, I, Acta Math. 127 (1971), 79–183, See also [1]. Hormander2

[8] , The analysis of linear partial differential operators, vol. 1, Springer-Verlag, Berlin,

Heidelberg, New York, Tokyo, 1983.

Munkres1

[9] James R. Munkres, Analysis on manifolds, Addison-Wesley Publishing Company Advanced

Book Program, Redwood City, CA, 1991.

Seeley1

[10] R.T. Seeley, Complex powers of elliptic operators, Proc. Symp. Pure Math. 10 (1967), 288–

307.

Spivak1

[11] Michael Spivak, Calculus on manifolds. A modern approach to classical theorems of advanced

calculus, W. A. Benjamin, Inc., New York-Amsterdam, 1965.

Wodzicki7

[12] M. Wodzicki, Noncommutative residue. I. Fundamentals, K-theory, arithmetic and geometry

(Moscow, 1984–1986), Lecture Notes in Math., Springer, Berlin-New York, 1987, pp. 320–399.

Wodzicki1

[13] , Cyclic homology of pseudodifferential operators and noncommutative Euler class, C. R. Acad. Sci. Paris 306 (1988), 321–325.